OpenStudy (anonymous):

Simplify using your knowledge of properties. [(1/3)(3)](2 + -1 + 1)

OpenStudy (anonymous):

The answer is y = 3. Collect the like terms together 6y = 10 + 8. This will be 6y = 18. Find the value of y by dividing both sides by 6. 6y/6 = 18/6.

OpenStudy (anonymous):

THANK YOU CUTIE :*

OpenStudy (anonymous):

u welcome